OpenStudy (anonymous):

Solve the system. then classify the system as consistent and independent, consistent and dependent or inconsistent. 2x+y=9 x+y=4 please work it out :)

OpenStudy (anonymous):

2x+y=9 --------1 x+y=4 ---------2 subtract 2nd equation from 1st equation 2x+y=9 -x-y=-4 ---------- x=5 now plug x value into one of the equations 5+y=4 y=4-5 y=-1 so,x=5 and y=-1 consistent and independent